# Second-Generation Planet Formation in Post-AGB Discs: Testing the Role of Gravitational Instability

### v1

@@ 4 Gravitational Instability in Typical Post-AGB Discs
Having used our GI framework on well-studied YSO systems, we now apply the same methodology to the post-AGB binary population.
Estimates of post-AGB disc gas masses, based on ALMA observations of molecular line emission (e.g. CO), typically fall in the range of $10^{-4}$$10^{-2}$ M (Bujarrabal_2013; Gallardo_Cava_2021, e.g.). In contrast, disc dust masses inferred from radiative transfer modelling of near- and mid-infrared interferometric data and spectral energy distributions (SEDs) can be significantly higher—reaching up to 0.10.10.20.2 M in a few individual cases (Kluska_2018; Corporaal_2023, e.g.). These estimates are sensitive to assumptions about grain properties, vertical structure, and disc inclination, and are often subject to model degeneracies. Interpreting such high dust masses with a canonical gas-to-dust ratio of $\sim$100  [^8] A key uncertainty in interpreting both gas and dust measurements is the gas-to-dust ratio in evolved systems. While a ratio of $\sim$100 is commonly assumed in protoplanetary discs, post-AGB discs may have undergone significant chemical evolution or gas depletion. The true ratio remains unconstrained, and thus dust-based masses may overestimate, and gas-based masses may underestimate, the total disc mass. would imply total disc masses of several solar masses, in tension with dynamical constraints and gas-based mass estimates Hillen_2015; Kluska_2018. To account for the uncertainties in total disc mass estimates—arising from the discrepancy between gas and dust measurements and the unknown gas-to-dust ratio—we adopt a conservative range of 0.010.010.10.1 M for the total (gas + dust) disc mass. We assess gravitational stability across this range using a power-law surface density profile, consistent with standard characterisations of post-AGB discs from observations (Section 2.2). Typical truncation radii lie between 100 and 500 au; for our fiducial case, we adopt $r_{\rm out}=100$ au to represent the more compact end of the distribution, where surface densities are highest and instability most likely.
We assume a central binary mass of $M_{1}+M_{2}=1.5$ M, consistent with typical post-AGB binaries (Oomen_2018, e.g.), and adopt $R_{\rm star}=100$ R and $T_{\rm star}=5,000$ K for the post-AGB star. Assuming a binary separation of $1$ au, we set the disc inner radius to the expected dynamical truncation radius of $r_{\rm in}=3$ au following Artymowicz1994.
@@
Although power-law indices with $n>2$ appear in some observational fits to post-AGB discs Kluska_2022, such profiles are not physically plausible for a self-gravitating disc without an inner cavity, as the enclosed mass would diverge. To reflect the observed structure and ensure physical consistency, we adopt a power-law (Equation 10) to parametrise the surface densities of post-AGB disks following how they are typically modelled in literature, and we also include an inner cavity in our analysis (Section 2.2, Equation 12) to avoid this divergence and to remain consistent with the typical observed structure of post-AGB discs. This is in contrast with how we have parametrised YSO disks (Equation 13) in the preceding section.
@@ 4.1 Testing the Stability of the Massive Disc in IRAS 08544-4431
Figure 4 shows the resulting Toomre profiles for both disc mass estimates. If the higher mass inferred by Corporaal_2023 is accurate, the disc reaches $Q<4$ in its outer regions—approaching but not exceeding the fragmentation threshold. This places IRAS 08544-4431 among the most gravitationally unstable post-AGB discs in our sample, though it remains above the threshold for fragmentation. By contrast, the lower mass estimate from Bujarrabal_2018 yields a Toomre profile that stays well within the stable regime across all radii. These results highlight the sensitivity of the Toomre parameter to disc mass assumptions and underscore the need for more robust constraints on gas and dust masses in evolved systems.
@@
In all cases, we adopt a power-law surface density profile (Equation 10) with $n=1$, and additionally show $n=0.5$ for post-AGB systems. For steep profiles ($n=3$), where $Q$ decreases with radius, we evaluate $Q$ at the inner radius ($r_{\rm in}$) instead of $r_{\rm out}$, as shown in the right panel of Figure 5. We adopt fixed parameters of $r_{\rm in}=3$ AU (corresponding to the dynamical truncation radius for a $T_{\rm subl}=1200$ K). The contours represent the value of $Q$ at the disc’s outer edge ($r=r_{\rm out}$), which corresponds to the minimum $Q$ for surface density slopes $n<2$. For $n>2$, where $Q$ decreases outward, the minimum occurs at the inner edge, and we report that value instead.
We find that only sufficiently massive discs are prone to enter the Toomre-unstable regime. Discs with representative post-AGB masses ($\sim 0.01$ M) are unlikely to becomes unstable under present-day conditions and support planet formation via gravitational instability. This is in contrast with discs around main-sequence stars and white dwarfs (Figure 6) which can become gravitationally unstable with much lower disc masses.
@@ 5 Were Discs Less Stable Earlier in Their Lifetimes?
More massive and radially compact discs are more prone to gravitational instability. If the discs we observe today were more massive and compact in the past, they may have been unstable at earlier epochs – even if they appear stable today. Here we calculate the viscous evolution of disks to determine their properties in the past and to determine whether, at that time, they may have been gravitationally unstable.
@@
In Figure 7, we plot the Toomre parameter $Q$ as a function of radius for several YSO discs and a representative post-AGB disc ($M_{\rm disc,now}=0.02$ M, $r_{\rm ch,now}=100$ au, and assuming $\gamma=1$), evolved backwards in time for different numbers of elapsed viscous timescales ($T=t/t_{\rm visc}+1$). We assume constant inner radii during viscous evolution. For post-AGB disks, the inner radius is set by the dynamical truncation radius, while for YSOs, it is set by the sublimation radius (taken here as the radius where $T_{\rm subl}=1200$ K). These boundaries depend on stellar properties rather than viscous evolution. Figure 7 shows that even at earlier epochs, some of the discs were likely gravitationally stable —particularly in the cases of PDS70, HD169142, and the post-AGB example. However, in the case of GQ Lup, HL Tau, and Elias 2-27, they could have been gravitationally unstable in their pasts.
Our analysis includes all YSO systems discussed in Section 3, except AB Aurigae and L1448 IRS3B, whose density profiles are not compatible with the form in Equation 13. For post-AGB discs, we assume they follow the viscously evolving form of Equation 13, thus assigning them a characteristic radius. This differs from the simpler power-law model used in earlier sections.
@@
To calculate these parameters, we adopt representative post-AGB values similar to that used in Section 4 and Figure 3. Disc ages for various assumed viscosity parameters ($\alpha$) are presented in Table 2. Importantly, varying $T=t/t_{\rm visc}+1$ alters both the inferred disc age and the initial characteristic radius $r_{\rm ch,0}$, but we find that the total system age remains relatively stable for each specific value of $\alpha$.
We find that the disc age does not vary significantly with changes in the scaled time $T$ for a given $\alpha$, since the viscous timescale itself adjusts accordingly. However, for fixed $T$, the disc age scales linearly with $\alpha$ (for $\gamma=1$). The stellar ages compiled from the literature are included for comparison and should be considered upper limits to the corresponding disc ages. In general, lower $\alpha$ values yield disc ages that exceed the stellar ages—consistent with the findings of 1998hartmann, who argued that $\alpha\approx 0.01$ is typical for YSO discs. For Elias 2-27 specifically, 2009Andrews derived $\alpha=0.03$ from accretion rate measurements, which supports our finding that an assumed $\alpha=0.01$ overestimates the disc age relative to the stellar age.
For post-AGB systems, the inferred ages are consistent with their expected lifetimes of $\sim$$10^{5}$ yr Bujarrabal_2016.
@@
  • The estimated ages for the central stars of these discs are 1 Myrs for HL Tauri Booth_2020, 1 Myrs for Elias 2-27 2009Andrews, 5.4 Myrs for PDS 70 Keppler_2019, 6 Myrs for HD 169142Hammond_2023, and 2-5 Myrs for GQ Lupi Wu_2017.
@@ 6 Planet Formation in Post-Common Envelope Discs
In this section we apply our gravitational instability framework to the post-common envelope binary NN Serpentis, where the presence of two proposed circumbinary planets has been interpreted as evidence for second generation planet formation in a remnant disc. These planets were first inferred from eclipse timing variations Beuermann_2010. Given their proximity to the central binary, it has been argued that they could not have formed during the system’s main sequence evolution and subsequently survived both the giant phase and the common-envelope expansion. Instead, it has been proposed that they formed within acircumbinary disc created during the common-envelope interaction.
@@
The existence of these circumbinary planets, however, has been the subject of debate. While the planets were first inferred from eclipse timing variationsBeuermann_2010; Beuermann_2013, and later studies assessed the dynamical stability of the proposed orbits, suggesting that such configurations could be stable 2012Horner; Marsh_2013, their interpretation as planets is regarded with some caution. In particular, Marsh_2013 noted that the planets are only indirectly detected, and that alternative mechanisms – most notably magnetic activity cycles (Applegate’s mechanism) – can also reproduce the observed eclipse timing variations.
@@
On the premise of the planets’ existence, 2014schleicher developed an analytical framework to estimate the properties of a circumbinary disk formed through common-envelope ejection. By applying gravitational instability theory, they demonstrated that the planetary masses and orbital separations inferred for NN Ser are consistent with formation in such an environment.
We apply our updated gravitational instability model to determine whether such a disk is indeed prone to making planets. We carry out our analysis for a disk mass like the one used by 2014schleicher, and for a disk mass that is 10 times smaller. The reason for this is that while some circumbinary gas has been detected around some post-CE binaries Li2025, recent theoretical developments have suggested that full envelope ejection is likely and less massive disks would then form by fall-back of a small amount of bound material 2015ivanova; 2020Reichardt.
In Figure 9 we show the radial variation of the Toomre parameter $Q$ for the two assumed disk masses, 0.1 M and 0.01 M. The central source is a white dwarf with with $M_{c}=0.535$ M, $R_{\rm star}=0.0189$ R, and $T_{\rm star}=57\,000$ K 2004Haefner. We have also assumed an outer circumbinary disk radius of $r_{\rm out}=100$ au (similar to Figure 3, with the only difference being that the central object is a white dwarf instead of a post-AGB star). As expected, more massive discs are more prone to gravitational instability. Notably, despite its high temperature, the compact size of the white dwarf results in a lower overall luminosity and, consequently, a cooler mid-plane disk temperature, compared to systems with larger, even if cooler, post-AGB stars. This reduced thermal support lowers the sound speed, thereby decreasing $Q$ and making the disc more susceptible to fragmentation. However, even with these more favourable conditions, only the more massive disk leads to possible planet formation by gravitational instability.
While 2014schleicher assumed that the entire disc around NN Ser is marginally unstable ($Q(r)\sim 1$), we now investigate whether this assumption holds when a more realistic temperature profile than the one they assumed is adopted. To do so, we use Equation 7, incorporating the observed properties of the central white dwarf as listed above. Following the framework proposed by 2014schleicher, we find an outer radius of $r_{\rm out}=7.50$ au for the disc, which is slightly different from the value they found ($r_{\rm out}=6.48$ au – this difference is further discussed in B). We also use their values of $M_{\rm disk}=0.146$M, a power-law index of $n=1$ for the surface density profile (Equation 10), and a mass of the central object of $M_{\rm core}+M_{2}=0.646$ M (see Table 3).
@@
Figure 10 shows the radial variation of the Toomre parameter, $Q$, for NN Ser using their proposed disk mass, a disk radius of $r_{\rm out}=7.50$ au and the temperature profile of Equation 7. The figure shows that even with a more realistic thermal structure, the disc becomes gravitationally unstable beyond $\sim 3$ au - precisely where the proposed planets are located. This analysis therefore confirms that the gravitational instability mechanism remains viable under more physically motivated conditions than those selected by 2014schleicher, but only if the disk mass is on the high, rather than low side (0.1 Mvs. 0.01 M (for further details see B).
Table 3: Parameters adopted or determined by 2014schleicher in their study of NN Ser. For details see B.
[…]
Disk Mass, $M_{\rm disk}$ 0.146 M
Ejected Mass, $M_{\rm ej}$ 1.24 M
Primary mass, $M_{1}$ 2 M
Companion mass, $M_{2}$ 0.111 M
Primary core mass, $M_{\rm c}$ 0.535 M
Primary radius, $R_{*}$ 185 R
Disc’s surface density index, $n$ 1
Disc’s outer radius, $r_{\rm out}$ 7.5 R [^11] The value we find for $r_{\rm out}$ is different from that of 2014schleicher, see B for more details.
Disc’s ang. mom., $L_{\rm dep}$ $1.2\times 10^{52}$ g cm s-1121212The value we find for $L_{\rm dep}$ is different from that of 2014schleicher, see B for more details.
Disc’s ang. mom. enhancement factor, $\alpha_{\rm L}$ 12.5
critical viscosity, $\alpha_{\rm crit}$ 0.3
Figure 9: The Toomre parameter $Q$ as a function of radius for discs with masses of 0.01 M (left panel) and 0.1 M (right panel) and parameters similar to Figure 3, but with a white dwarf as the central star ($M_{c}=0.535$ M, $R_{\rm star}=0.0189$ R, and $T_{\rm star}=57,000$ K), and an outer disc radius of $r_{\rm out}=100$ au.
@@
Figure 10: The Toomre parameter, $Q$, as a function of orbital radius for the progenitor disc of NN Serpentis, based on the values of 2014schleicher, which are summarised in Table 3. The white dwarf parameters were $T_{\rm eff}=57,000$ K and $R_{\rm star}=0.0189$ R, and the mass of the central binary was set to be $M_{\rm c}+M_{2}=0.646$ M. We note that the outer radius of the disc used to reproduce this plot was $r_{\rm out}=7.50$ au, which we recomputed using the value we find for angular momentum in B.

### v2

@@ 4 Gravitational Instability in Typical Post-AGB Discs
Having used our gravitational instability framework on well-studied YSO systems, we now apply the same methodology to the post-AGB binary population.
Estimates of post-AGB disc gas masses, based on ALMA observations of molecular line emission (e.g., CO), typically fall in the range of $10^{-4}$$10^{-2}$ M (Bujarrabal_2013; Gallardo_Cava_2021, e.g.). In contrast, disc dust masses inferred from radiative transfer modelling of near and mid-infrared interferometric data and spectral energy distributions (SEDs) imply significantly higher total disc masses - reaching up to 0.10.10.20.2 M in a few individual cases (Kluska_2018; Corporaal_2023, e.g.,). These estimates are sensitive to assumptions about grain properties, vertical structure, and disc inclination, and are often subject to model degeneracies. Interpreting such high dust masses with a canonical gas-to-dust ratio of $\sim$100  [^8] A key uncertainty in interpreting both gas and dust measurements is the gas-to-dust ratio in evolved systems. While a ratio of $\sim$100 is commonly assumed in protoplanetary discs, post-AGB discs may have undergone significant chemical evolution and/or gas depletion. The true ratio remains unconstrained, and thus dust-based masses may overestimate, and gas-based masses may underestimate, the total disc mass. would imply total disc masses of order $10^{-1}$ M, in tension with dynamical constraints and gas-based mass estimates Hillen_2015; Kluska_2018. To account for the uncertainties in total disc mass estimates - arising from the discrepancy between gas and dust measurements and the unknown gas-to-dust ratio - we adopt a conservative range of 0.010.010.10.1 M for the total (gas + dust) disc mass. We assess gravitational stability across this range using a power-law surface density profile, consistent with standard characterisations of post-AGB discs from observations (Section 2.2). Typical truncation radii lie between 100 and 500 au; for our fiducial case, we adopt $r_{\rm out}=100$ au to represent the more compact end of the distribution, where surface densities are highest and instability most likely.
We assume a central binary mass of $M_{1}+M_{2}=1.5$ M, consistent with typical post-AGB binaries (Oomen_2018, e.g.,), and adopt $R_{\rm star}=100$ R and $T_{\rm star}=5,000$ K for the post-AGB star. Assuming a binary separation of $1$ au, we set the disc inner radius to the expected dynamical truncation radius of $r_{\rm in}=3$ au following Artymowicz1994.
@@
To reflect the observed structure and ensure physical consistency, we adopt a power-law (Equation 10) to parametrise the surface densities of post-AGB discs following how they are typically modelled in literature, and we also include an inner cavity in our analysis (Section 2.2, Equation 12) to avoid this divergence and to remain consistent with the typical observed structure of post-AGB discs. This is in contrast with how we have parametrised YSO discs (Equation 13) in the preceding section.
@@ 4.1 Testing the Stability of the Massive Disc in IRAS 08544-4431
Figure 4 shows the resulting Toomre profiles for both disc mass estimates. If the higher mass inferred by Corporaal_2023 is accurate, the disc reaches $Q<4$ in its outer regions - approaching but not exceeding the fragmentation threshold. This places IRAS 08544-4431 among the most gravitationally unstable post-AGB discs in our sample, though it remains above the threshold for fragmentation. By contrast, the lower mass estimate from Bujarrabal_2018 yields a Toomre profile that stays well within the stable regime across all radii. These results highlight the sensitivity of the Toomre parameter to disc mass assumptions and underscore the need for more robust constraints on gas and dust masses in evolved systems.
@@
In all cases, we adopt a power-law surface density profile (Equation 10) with $n=1$, and additionally show $n=0.5$ for post-AGB systems. For steep profiles ($n=3$), where $Q$ decreases with radius, we evaluate $Q$ at the inner radius ($r_{\rm in}$) instead of $r_{\rm out}$, as shown in the right panel of Figure 5. We adopt fixed parameters of $r_{\rm in}=3$ au (corresponding to the dynamical truncation radius for a post-AGB binary system with an orbital separation of $a=1$ au). The contours represent the value of $Q$ at the disc’s outer edge ($r=r_{\rm out}$), which corresponds to the minimum $Q$ for surface density slopes $n<2$. For $n>2$, where $Q$ decreases outward, the minimum occurs at the inner edge, and we report that value instead.
We find that only sufficiently massive and compact discs are prone to enter the Toomre-unstable regime. Discs with representative post-AGB disc masses ($\sim 0.01$ M) are unlikely to becomes unstable under present-day conditions and support planet formation via gravitational instability. This is in contrast with discs around main-sequence stars and white dwarfs (Figure 6) which can become gravitationally unstable with much lower disc masses.
@@ 5 Were Discs Less Stable Earlier in Their Lifetimes?
More massive and radially compact discs are more prone to gravitational instability. If the discs we observe today were more massive and compact in the past, they may have been unstable at earlier epochs – even if they appear stable today. Here we use viscous evolution theory of discs to determine their properties in the past and to determine whether, at that time, they may have been gravitationally unstable.
@@
In Figure 7, we plot the Toomre parameter $Q$ as a function of radius for several YSO discs and a representative post-AGB disc ($M_{\rm disc,now}=0.02$ M, $r_{\rm ch,now}=100$ au, and assuming $\gamma=1$), evolved backwards in time for different numbers of elapsed viscous timescales ($T=t/t_{\rm visc}+1$). We assume constant inner radii during viscous evolution. For post-AGB discs, the inner radius is set by the dynamical truncation radius, while for YSOs, it is set by the sublimation radius (taken here as the radius where $T_{\rm subl}=1200$ K). These boundaries depend on stellar properties rather than viscous evolution. Figure 7 shows that even at earlier epochs, some of the discs were likely gravitationally stable - particularly in the cases of PDS 70, HD169142, and the post-AGB example. However, in the case of GQ Lup, HL Tau, and Elias 2-27, they could have been gravitationally unstable in their pasts.
Our analysis includes all YSO systems discussed in Section 3, except AB Aurigae and L1448 IRS3B, whose previously estimated density profiles are not compatible with the form in Equation 13. For post-AGB discs, we assume they follow the viscously evolving form of Equation 13, thus assigning them a characteristic radius. This differs from the simpler power-law model used in earlier sections.
@@
To calculate these parameters, we adopt representative post-AGB values similar to those used in Section 4 and Figure 3. Disc ages for various assumed viscosity parameters ($\alpha$) are presented in Table 2. Importantly, varying $T=t/t_{\rm visc}+1$ alters both the inferred disc age and the initial characteristic radius $r_{\rm ch,0}$, but we find that the total system age remains relatively stable for each specific value of $\alpha$.
We find that the disc age does not vary significantly with changes in the scaled time $T$ for a given $\alpha$, since the viscous timescale itself adjusts accordingly. However, for fixed $T$, the disc age is inversely proportional to $\alpha$ for $\gamma=1$. The stellar ages compiled from the literature are included for comparison and should be considered upper limits to the corresponding disc ages. In general, lower $\alpha$ values yield disc ages that exceed the stellar ages - consistent with the findings of 1998hartmann, who argued that $\alpha\approx 0.01$ is typical for YSO discs. For Elias 2-27 specifically, 2009Andrews derived $\alpha=0.03$ from accretion rate measurements, which supports our finding that an assumed $\alpha=0.01$ overestimates the disc age relative to the stellar age.
For post-AGB systems, the inferred ages are consistent with their expected lifetimes of $\sim$$10^{5}$ yr Bujarrabal_2016 for the higher end of viscosity parameters we adopted; $\alpha=0.01$.
@@
  • The estimated ages for the central stars of these discs are 1 Myr for HL Tauri Booth_2020, 1 Myr for Elias 2-27 2009Andrews, 5.4 Myr for PDS 70 Keppler_2019, 6 Myr for HD 169142Hammond_2023, and 2-5 Myr for GQ Lupi Wu_2017.
@@ 6 Planet Formation in Post-Common Envelope Discs
In this section we apply our gravitational instability framework to the post common envelope binary NN Ser, where the presence of two proposed circumbinary planets has been interpreted as evidence for second generation planet formation in a remnant disc. These planets were first inferred from eclipse timing variations Beuermann_2010. Given their proximity to the central binary, it has been argued that they could not have formed during the system’s main sequence evolution and subsequently survived both the giant phase and the common envelope expansion. Instead, it has been proposed that they formed within a circumbinary disc created during the common envelope interaction.
@@
The existence of these circumbinary planets, however, has been the subject of debate. While the planets were first inferred from eclipse timing variations Beuermann_2010; Beuermann_2013, and later studies concluded the proposed orbits to be dynamically stable 2012Horner; Marsh_2013, the interpretation of the observations as being due to planets is nonetheless regarded with caution. In particular, Marsh_2013 noted that the planets are only indirectly detected, and that alternative mechanisms – most notably magnetic activity cycles (Applegate’s mechanism 1992Applegate) – can also reproduce the observed eclipse timing variations.
@@
On the assumption of the planets’ existence, 2014schleicher developed an analytical framework to estimate the properties of a (non observed) circumbinary disc formed by a common envelope ejection. Then applying gravitational instability theory, they demonstrated that the planetary masses and orbital separations inferred for NN Ser are consistent with planet formation in such an environment. Here we connect our gravitational instability framework applied to (observed) post-AGB discs to their analysis of NN Ser, and in so doing we re-examine several aspects of their model.
By considering common envelope physics, 2014schleicher derived a disc mass of $M_{\rm disk}=0.146$ M [^11] It is not clear what post-CE disc masses may be, with simulations indicating the ability of the interaction to eject almost the entire envelope leaving very little mass to form a disc (2015ivanova; 2020Reichardt; Gonzalez2022, e.g.,). Observationally, only very low mass discs have been detected around some post-CE binaries Li2025.. With this disc mass, they went on to assume that the entire disc around NN Ser is marginally gravitationally unstable ($Q(r)\sim 1$) in order to reproduce the observed planet masses.
Here we use our framework to determine whether their disc would be unstable to planet formation. We use their disc mass, their surface density (a power-law index of $n=1$ for the surface density profile – Equation 10), and their central binary mass, $M_{\rm core}+M_{2}=0.646$ M. However, we use a more realistic temperature profile than used by them, in line with our analysis - Equation 7, where the disc is heated by irradiation from the central source, which in the case of NN Ser is a white dwarf. Finally, we use a disc outer radius of $r_{\rm out}=7.50$ au, calculated by conservation of angular momentum before and after disc formation - see C for details.
@@
Figure 9 shows the radial variation of the Toomre parameter, $Q$, for NN Ser. The figure shows that even with a more realistic thermal structure, the disc becomes gravitationally unstable beyond $\sim 3$ au - precisely where the proposed planets are located. This analysis therefore confirms that the gravitational instability mechanism remains viable under more physically motivated conditions than those selected by 2014schleicher, but only if the disc mass is as high as they propose.
@@
Figure 9: The Toomre parameter, $Q$, as a function of orbital radius for the progenitor disc of NN Ser, based on the values of 2014schleicher, which are summarised in Table 3. The white dwarf parameters were $T_{\rm eff}=57,000$ K and $R_{\rm star}=0.0189$ R, and the mass of the central binary was set to be $M_{\rm c}+M_{2}=0.646$ M. We note that the outer radius of the disc used to reproduce this plot was $r_{\rm out}=7.50$ au, which we recomputed using the value we find for angular momentum in C.